Determine whether the vectors u and v are parallel, orthogonal, or neither.

u = <6, -2>, v = <8, 24> 


1
Expert's answer
2021-01-25T00:13:15-0500

Two vectors u and v are orthogonal if dot product u•v=0 .

u•v=6•8+(-2)•24=48-48=0"\implies"vectors u and v are orthogonal.

Answer: u and v are orthogonal.